# Bullet: Boosting GPU Utilization for LLM Serving via Dynamic Spatial-Temporal Orchestration

### v1

@@ 3.5.1. Execution Workflow
Figure 9 shows the concurrent execution engine for prefill (green boxed) and decode (orange), isolated in separate processes and operating independently with decentralized schedulers. We exemplify a scenario in which the prefill engine computes the last groups of layers for request 5 while the decode engine executes requests 1-4. Taking the control and dataflow for request 5 as an example, the execution begins with the scheduling workflow depicted in section 3.3.1, illustrated as the purple box in the engine. Initially, in time step 1, system status from the decode engine is ❶ retrieved from the CPU metadata buffer, enabling the prefill engine to perceive holistic information for scheduling, including decode batch size, real-time progress and latencies. Conversely, the request metadata, such as KV cache indices (Zheng et al., 2023) and progress, are sent to the buffer to share with decode engine. After that, the SLO-aware scheduler determines the optimal resource configuration and ❷ launches kernels from the layer group to the designated SM partition. Note that the decode engine is performing the same process and executing decoding concurrently with the prefill. When the prefill kernels finish, migrating request 5 to the decode engine is copy-free, since the KV cache and index table are shared by the two engines with GPU memory pool (section 3.5.2). In time step 2, when the decode engine finishes an iteration and starts to schedule, the engine ❸ fetches system status from the buffer, similar to the process in ❶. Therefore, request 5’s metadata is received as a component of the status by the decode engine, allowing for the scheduler to batch request 5 with existing decoding requests. Finally, the scheduler follows the same workflow for ❹ resource determination and kernel launch, with prefill engine computing on request 6 concurrently.
Bullet’s scheduler works independently while proactively communicating through the shared metadata buffer, unblocking both CPU and GPU execution. This decentralized architecture allows kernels to submit to GPU while eliminate the need for frequent synchronization compared with a centralized controller.
3.5.2. Memory Space Sharing
To facilitate inter-engine communication, Bullet designs a shared CPU buffer and unified GPU memory pool. The CPU buffer is implemented as OS-managed shared memory, storing system status and request metadata. During runtime initialization, the prefill engine creates the shared memory, and the decode engine maps to the region. This allows engines to perform direct read/write operations on the buffer with low-latency. Bullet also defines several control bits in the buffer to indicate the data availability. For GPU memory management, Bullet employs a dedicated initialization process that allocates model weights and KV cache (Kwon et al., 2023) prior to engine launch. The GPU memory region is exported as a handle using CUDA’s cudaIpcGetMemHandle API (Corporation, 2025), which is then transmitted to the engines via ZeroMQ (ZeroMQ authors, [n. d.]). This process is lightweight and causes no adverse effects on the memory allocation, as documented. Upon receiving the handle, each engine maps the memory using cudaIpcOpenMemHandle, gaining direct access to model weights and KV cache within its own memory space without performance penalty, which is similar to OS-managed shared memory.
ShareGPT | Azure-Code | arXiv-Sum | |
norm. TTFT (ms) | 3.0 | 1.5 | 1.5 |
TPOT (ms) | 150 | 200 | 175 |
@@ 4.1. Methodology
Platforms. We conduct experiments on a server equipped with Nvidia A100-PCIe-80GB GPUs, featuring 108 SMs and 2TB/s of HBM bandwidth, with GPU clocks locked to the maximum stable frequency of 1410MHz. The CPU is Intel Xeon Gold 6150 CPU, 256GB of DRAM, and runs Debian Linux 5.10.0-33. The version of the Nvidia driver is 535.183.06 and CUDA Toolkit is 12.2.
Implementation. We implement Bullet on top of SGLang (Zheng et al., 2023) v0.3.0 and PyTorch 2.4.0 with 4100 lines of Python code, and integrate a modified SM partitioning technique (Bakita and Anderson, 2023, 2024) to optimize GPU resource allocation within the serving engine. The prefill and decode engines are implemented as SGLang’s workers launched in separate processes, with MPS (NVIDIA, 2025) enabled for spatial sharing. The end-to-end processing pipeline begins with SGLang’s built-in tokenizer handling incoming requests. Tokenized requests are routed to the prefill engine with metadata pushed to the shared CPU buffer. Then the standard execution workflow (section 3.5) starts, with generated tokens passing through SGLang’s native detokenizer, and finally delivers to the front-end server.
Evaluated Schemes. We compare Bullet against several state-of-the-arts implementations of chunked prefill-based optimization: vLLM (Kwon et al., 2023) V1 (chunk size: 1024), SGLang (Zheng et al., 2023) v0.3.0 (chunk sizes: 1024/2048), and Nanoflow (Zhu et al., 2024) (chunk size: 1024).
Models and Datasets. We evaluate the system using the Llama-3.1-8B model (Dubey et al., 2024) across three representative datasets, which are widely used in previous works (Kwon et al., 2023; Agrawal et al., 2023a; Zhu et al., 2024), and the arrival distribution of requests also follows a Poisson process. Figure 10 details the cumulative distribution function (CDF) of the datasets.The ShareGPT (Teams., 2023) contains real-world conversational data, Azure-Code (Patel et al., 2024) is production code completions by Azure, and arXiv-Summary (Cohan et al., 2018a, b) is long-context summarization.
Metrics We report TTFT, TPOT and throughput aligned with previous works (Kwon et al., 2023; Agrawal et al., 2023a; Zhu et al., 2024). For SLO compliance (goodput) measurement, we define as prefill and decode latency satisfy both the constraints, as listed in Table 2. Due to the variation in sequence length, we use normalized input latency (norm. TTFT) established in previous work (Wu et al., 2024). We collect hardware metrics using Nvidia Nsight Systems 2023.2.3 (nsy, 2025), a low-overhead profiling tool.
@@ 4.2. End-to-End Performance
We evaluate Bullet against four baselines across three production workloads with varying request rates (Figure 11). Bullet demonstrates consistent throughput improvements, achieving 1.09 average and up to 1.20 higher throughput compared to SGLang-1024. Overall, Bullet maintains superior prefill latency enhancement (13.5) with acceptable TPOT (0.94) in average through dynamic SM allocation, translating to 1.86 end-to-end speedup compared with SGLang-1024. In contrast, all the chunked prefill-based systems exhibit unacceptable TTFT degradation even in a low request rate. For example, on the ShareGPT workload, with request rate of 20req/s, the prefill latency ranges from 1.8s (Nanoflow) to 14.9s (vLLM). The key reason is that chunk prefill limits the system’s capability to process the prefill phase (section 2.3.2) and creates a cascading congestion for pending requests. By eliminating these bottlenecks through intelligent SM allocation and adaptive scheduling, Bullet achieves mean TTFT of 0.16s and P90 tail latency of 0.31s, which is 54.9 and 78.5 better to SGLang-1024. The lower prefill latency remarkably translates to larger decode batches that boost overall throughput and SLO compliance (1.49).
Bullet effectively balances the throughput-latency tradeoff that skewed in chunked prefill systems. While SGLang-2048 demonstrates a 3.20 TTFT improvement and 1.18 higher throughput than SGLang-1024, the 2048 chunk suffers from 0.78 worse TPOT in average, confirming the imbalanced tradeoff inherent in chunked prefill. Bullet breaks this paradigm by simultaneously achieving both lower TTFT (4.2) and better TPOT (1.20) than SGLang-2048 through exploiting GPU utilization via concurrent prefill-decode execution. While Nanoflow achieves 2.37 longer TTFT and 0.86 shorter TPOT than Bullet via static kernel overlapping pipeline, it still suffers from the inherent limitations of chunked prefill that inflate TTFT tail latency. This results in 5.2% lower SLO compliance and 20.0% lower throughput compared to Bullet.
@@ 4.3.1. Dynamic SM Partition
We break down the effectiveness of dynamic resource provisioning in Figure 12 by timeline with the Azure-Code workload (5.0 req/s). The top row in Figure 12(a) demonstrates the number of SMs provisioned for the prefill phase, with each bar showing the SM count and duration. On request rate bursts (spikes in Figure 12-bottom), Bullet adaptively sets the number of prefill SMs to use full GPU, and may temporarily delay decode requests. This enables the rapid response for queuing requests, avoiding excessive waiting time. After processing the pending requests, Bullet quickly reconfigurates the resources to a balance point for both phases. Instead, chunked prefill-based systems are not able to optimize such a scenario, since redundant KV cache access inflates the prefill speed. Moreover, decode batch shares the token budget with prefill tokens (Figure 12(b)), enforcing more iterations to complete a prefill, which further degrades performance. These factors jointly result in 4.17 longer queuing delay for SGLang-2048. Bullet exempts from the token budget (Figure 12-middle) and orchestrates the two phases with fine-grained resource control to saturate GPU, significantly decreasing both TTFT and TPOT by 9.15 and 1.33, respectively.
@@ 4.3.2. Overheads
We quantify the CPU overhead of Bullet’s key components in Table 3 Sending and receiving metadata only necessitates serialization and de-serialization of Python objects, therefore achieving mean latency of 0.21. The performance prediction module incurs merely 10.2 overhead by simply invoking the analytical model. For GPU resource management, pre-allocated CUDA streams with configured SM partitions enable instantaneous reconfiguration, maintaining negligible runtime overhead. These design choices jointly ensure Bullet’s control plane adds minimal latency while enabling fine-grained resource optimization.
@@ 4.5. Ablation Studies
4.5.1. Effect of different components
Figure 14 analyzes the contributions of different components in Bullet. Three comparison systems isolate part of the design. Naive: concurrent execution without resource provisioning or scheduling. w/Partition: adds resource provision only. w/Scheduler: Incorporates only request reordering and delayed decode. The Naive design exhibits the expected latency imbalance (section 4.4), high TPOT and low TTFT attributed to unpartitioned resource contention. The w/Partition variant improves TPOT for Azure-Code but suffers unacceptable TTFT degradation from its inability to reorder pending requests. Conversely, w/Scheduler maintains comparable latency while reducing Azure-Code TPOT through contention alleviation. Only with the complete Bullet design, incorporating both partitioning and scheduling, achieve balanced latency across all workloads.
4.5.2. Performance Estimator Accuracy
We evaluate the precision of the profile-augmented analytical model in Figure 15. On average, the model achieves 88% accuracy in SLO compliance estimation, i.e., classifying whether requests will meet or violate latency targets. While the mean relative error between predicted and actual durations is 19.1%, this absolute error is inconsequential for scheduling decisions. The scheduler only requires reliable identification of SLO violations, which our model delivers through its high compliance prediction accuracy. This confirms that the model is fully sufficient for guiding Bullet’s real-time scheduling.
@@ 5. Related Works
LLM Serving Systems. Optimizations for LLM serving have been extensively explored from multiple system layers. At the kernel level, current systems heavily rely on high-performance operator libraries, especially hardware vendor-optimized GEMM implementations (NVIDIA Corporation, 2025; AMD, 2025; Osama et al., 2023) and specialized attention kernels (Dao et al., 2022; Sanovar et al., 2024; Prabhu et al., 2025) . While achieving near-peak performance for individual operators, the dynamic computational demands of end-to-end LLM serving create system-level utilization gaps. Bullet builds upon these kernel optimizations while introducing adaptive phase coordination to address these gaps. Scheduling innovations like continuous batching (Yu et al., 2022) and chunked prefill (Agrawal et al., 2023a) improve batching efficiency but maintain lockstep prefill and decode phase execution that underutilizes hardware. In contrast, Bullet fully parallelizes the phases to maximize GPU saturation. This approach also enhances emerging intra-device parallelism techniques (Duan et al., 2024; Zhu et al., 2024), further validating its effectiveness. Recent works disaggregate prefill and decode phases on separate replicas (Zhong et al., 2024; Patel et al., 2024; Hu et al., 2024; Gao et al., 2024; Qin et al., 2025; DeepSeek-AI et al., 2024) to eliminate the interference, but require costly KV cache migration and workload-specific GPU allocation, limiting the implementation in high bandwidth interconnect clusters. As such, Bullet is orthogonal to these works and offers complementary advantages in the transitional mixed instances required by disaggregated systems when switching between prefill and decode modes.
@@
Concurrent Kernel Execution. Co-executing kernels with complementary resource demands have been proven to improve GPU utilization and performance (Zhang et al., 2025a; Strati et al., 2024; Chen et al., 2022; Lin et al., 2023). Spatial-temporal multiplexing techniques (Strati et al., 2024; Choi et al., 2022; Chow et al., 2023; Dhakal et al., 2020; Gu et al., 2023; Chen et al., 2022; Kwon et al., 2020) typically classify kernels by computational characteristics and rely on CUDA streams (Corporation, 2025) or MPS (NVIDIA, 2025) for kernel submission. However, these approaches treat the hardware scheduler as a blackbox, providing limited control (Jain et al., 2019; Gilman et al., 2020; Olmedo et al., 2020) over actual execution. Alternative solutions employ static kernel fusion (Jangda et al., 2024; Li et al., 2022; Wang et al., 2010; Han et al., 2022; Wang et al., 2016) and explore optimization space with compiler (Ma et al., 2020; Hwang et al., 2023; Zhuang et al., 2024; Wu et al., 2025) to control the execution precisely. Nonetheless, the kernel-level modifications lack adaptability for online dynamic workloads. While some systems use precise kernel timing and resource provisioning (Zhang et al., 2025a; Han et al., 2024; Zhang et al., 2025b), they focus on traditional DNN workloads and fail to address LLM-specific requirements, including prefill-decode phase coordination, KV cache management and iteration-level batching. Bullet distinguishes itself from previous works by fine-grained orchestration of LLM’s phases and layer-level precise resource management.
@@
Task Scheduling. Concurrent kernel execution requires effective scheduling to enhance performance. Several works (Chen et al., 2022; Lin et al., 2023) schedule concurrent kernels based on inter-kernel dependencies, while others (Han et al., 2022; Shubha et al., 2024; Strati et al., 2024) classify tasks as real-time or best-effort and apply prioritized scheduling. Latency model-based solutions (Zhang et al., 2025a; Han et al., 2024; Ng et al., 2023; Wu et al., 2023; Zhang et al., 2025b) schedule kernels within latency constraints but rely on excessive offline profiling, which is unsuitable for the dynamic workloads in LLM serving. Although estimating standalone LLM inference is well-established (Agrawal et al., 2024; Yuan et al., 2024), the complex interactions between concurrent prefill and decode phases remain unstudied. Bullet addresses this gap by developing a profile-augmented analytical model for performance estimation and refining kernel scheduling mechanisms for LLM serving.

### v2

@@ 3.5.2. Execution Workflow
The prefill engine’s execution begins with receiving requests and follows the scheduling workflow in section 3.3, retrieving decode-side states from the buffer, such as batch size and TPOTs. Once prefill completes, the request migrates to decode without KV-cache transfer, and only metadata is asynchronously sent to the decode engine via ZeroMQ (ZeroMQ authors, [n. d.]), enabling microsecond overhead. The decode engine receives newly prefilled requests and merges them into the current running batch. Output tokens are directly forwarded to the frontend server, eliminating any CPU involvement by the prefill engine. Bullet’s control plane works independently while proactively communicating through the buffer, unblocking both CPU and GPU execution. This decentralized architecture allows concurrent kernel submissions while eliminating the need for frequent synchronization compared with a centralized architecture. We implement Bullet on top of SGLang (Zheng et al., 2023) v0.4.6 and PyTorch 2.6.0 with 4100 lines of Python code, and integrate a modified libsmctrl (Bakita and Anderson, 2023) library to optimize GPU resource allocation within the serving engine. The prefill and decode engines are implemented as SGLang’s workers, with MPS (NVIDIA, 2025) enabled for spatial sharing.
@@ 4.1. Methodology
Evaluated Schemes. We compare Bullet against several state-of-the-art systems of chunked prefill-based optimization: vLLM (Kwon et al., 2023) v0.8.5 (1024-chunk), SGLang (Zheng et al., 2023) v0.4.6 (1024/2048-chunk), and Nanoflow (Zhu et al., 2025) (1024-chunk).
Workloads Three representative datasets are used for evaluation, whose request arrivals follow a Poisson process, as widely adopted in prior works (Kwon et al., 2023; Agrawal et al., 2023; Zhu et al., 2025). Figure 13 details the cumulative distribution function (CDF) of the datasets’ sequence length. The ShareGPT (Teams., 2023) contains real-world conversational data, Azure-Code (Patel et al., 2024) is production code completion trace released by Azure, and arXiv-Summary (Cohan et al., 2018) is long-context summarization.
Models and Platforms. Experiments are conducted on three servers in Table 3 with CUDA 12.4. We evalute dense model Llama3.1-8B/70B (Dubey et al., 2024) and MoE model Qwen3-235B-A22B-FP8 (Yang et al., 2025). All multi-GPU setups use tensor parallelism.
GPUs | Network Bw. | #SMs/GPU | |
8A100-80GB | 20 GB/s | 108 | |
8H100 | 600 GB/s | 132 | |
8H20 | 400 GB/s | 78 |
ShareGPT | Azure-Code | arXiv-Sum | |
norm. TTFT | 3.0 ms | 1.5 ms | 1.5 ms |
TPOT | 150 ms | 200 ms | 175 ms |
Metrics. We report TTFT, TPOT and throughput. For SLO compliance measurement, we define as prefill and decode latency satisfying both the constraints listed in Table 4, aligned with previous works (Zhu et al., 2025; Wu et al., 2024; Zhong et al., 2024). Due to the variation in sequence length, we use normalized input/generation latency established in previous works (Wu et al., 2024; Zhu et al., 2025) for SLO measurement. We collect hardware metrics using Nsight Systems (nsy, 2025).
@@ 4.2. End-to-End Performance
4.2.1. Single GPU Performance
Figure 14 evaluate Bullet against four baselines across three real-world workloads. Bullet demonstrates consistent throughput improvements, achieving 1.09 average and up to 1.20 higher throughput compared to SGLang-1024. Overall, Bullet maintains superior prefill latency enhancement (13.5) with acceptable TPOT (0.94) in average through dynamic SM allocation, translating to 1.86 end-to-end speedup compared with SGLang-1024. In contrast, all the chunked prefill-based systems exhibit unacceptable TTFT degradation even in a low request rate. For example, on the ShareGPT workload, with request rate of 20req/s, the prefill latency ranges from 1.8s (Nanoflow) to 14.9s (vLLM). The key reason is that chunk prefill limits the system’s capability to process the prefill phase (section 2.3.2) and creates a cascading congestion for pending requests.
4.2.2. SLO Attainment
Figure 14(d) presents P90 SLO attainment rate. By eliminating these bottlenecks through intelligent SM allocation and adaptive scheduling, Bullet achieves mean TTFT of 0.16s and P90 tail latency of 0.31s, which is 54.9 and 78.5 better to SGLang-1024. The lower prefill latency remarkably translates to larger decode batches that boost overall throughput and SLO compliance (1.49). Bullet effectively balances the throughput-latency tradeoff that is skewed in chunked prefill systems. While SGLang-2048 demonstrates a 3.20 TTFT improvement and 1.18 higher throughput than SGLang-1024, the 2048 chunk suffers from 0.78 worse TPOT on average, confirming the imbalanced tradeoff inherent in chunked prefill. Bullet breaks this paradigm by simultaneously achieving both lower TTFT (4.2) and better TPOT (1.20) than SGLang-2048 through exploiting GPU utilization via concurrent prefill-decode execution. While Nanoflow achieves 2.37 longer TTFT and 0.86 shorter TPOT than Bullet via static kernel overlapping pipeline, it still suffers from the inherent limitations of chunked prefill that inflate TTFT tail latency. This results in 5.2% lower SLO compliance and 20.0% lower throughput compared to Bullet.
4.2.3. Multi-GPU Performance
Figure 15 depicts the generation latency of Bullet when Llama3.1-70B is deployed across eight A100 GPUs, with vLLM and SGLang using a chunk size of 2048. On the ShareGPT workload at 3.0 req/s, Bullet achieves a mean latency of 173 ms/token (223 ms/token for P99), whereas vLLM and SGLang attain 207 ms/token and 319 ms/token, respectively. For the Azure-Code dataset, where input lengths are significantly longer, Bullet’s advantages is more pronounced, sustaining 203 ms/token at 0.5 req/s. In contrast, vLLM and SGLang already exhibit 213 ms/token and 291 ms/token at a lower rate of 0.3 req/s.
On H100 GPUs, the elevated memory bandwidth mitigates chunked-prefill overhead, yet Bullet continues outperforms SGLang and vLLM. Bullet’s concurrent handling of prefill and decode tasks leverages potential overlaps between computation, communication and memory operations with fine-grained SM control between phases. We also demonstrate Bullet’s generalization to other LLMs. For the highly sparse, FP8-quantized MoE Qwen3-235B-A22B, chunked prefill’s inherent lock-step batching impedes both prefill and decode speeds. However, Bullet orchestrates the two phases effectively. At 4.0 req/s, Bullet delivers 110 ms/token, with 1.4s TTFT and 45ms TPOT—respectively, which is 17.7, 16.2 and 4.2 lower than vLLM.
@@ 4.3.1. Dynamic SM Partition
We break down the effectiveness of dynamic resource provisioning in Figure 17 by timeline with the Azure-Code workload (5.0 req/s, Llama3.1-8B). The top row in Figure 17(a) demonstrates the number of SMs provisioned for the prefill phase, with each bar showing the SM count and duration. On request rate bursts (spikes in Figure 17-bottom), Bullet adaptively sets the number of prefill SMs to use full GPU, and may temporarily delay decode requests. This enables the rapid response for queuing requests, avoiding excessive waiting time. After processing the pending requests, Bullet quickly reconfigurates the resources to a balance point for both phases. Instead, chunked prefill-based systems are not able to optimize such a scenario, since redundant KV cache access inflates the prefill speed. Moreover, decode batch shares the token budget with prefill tokens (Figure 17(b)), enforcing more iterations to complete a prefill, which further degrades performance. These factors jointly result in 4.17 longer queuing delay for SGLang-2048. Bullet exempts from the token budget (Figure 17-middle) and orchestrates the two phases with fine-grained resource control to saturate GPU, significantly decreasing both TTFT and TPOT by 9.15 and 1.33, respectively.
4.3.2. GPU Utilization
We measure hardware utilization with Nsight Systems and present it in Figure 18. Between 0 s and 27 s, when the system concurrently handles prefill and decode requests, Bullet sustains an average of 86.2% active SM cycles, which is 11.2% higher than SGLang. Correspondingly, utilization of Tensor Cores rises by 11.8%, while memory-bandwidth utilization increases by 19.3%. These utilization gains directly reduce the mean end-to-end request latency from 25.8 s to 21.4 s.
@@ 4.3.3. Overheads
Table 5 quantifies the CPU overhead of Bullet’s key components. Sending and receiving metadata only necessitates serialization and deserialization of Python objects, therefore achieving mean latency of 0.21. The performance prediction module incurs merely 10.2 overhead by simply invoking the analytical model. For GPU resource management, SM masking enables instantaneous reconfiguration with negligible runtime overhead. These designs jointly ensure Bullet’s control plane adds minimal latency while enabling fine-grained resource optimization.
@@ 4.5. Ablation Studies
Figure 20 analyzes the contributions of different components in Bullet by isolating part of the design. Naive: concurrent execution without resource provisioning or scheduling. w/Partition: adds resource provision only. w/Scheduler: Incorporates only request reordering and delayed decode. The Naive design exhibits the expected latency imbalance (section 4.4), high TPOT and low TTFT attributed to unpartitioned resource contention. The w/Partition variant improves TPOT for Azure-Code but suffers unacceptable TTFT degradation from its inability to reorder pending requests. Conversely, w/Scheduler maintains comparable latency while reducing Azure-Code TPOT through contention alleviation. Only the complete Bullet design, combining both partitioning and scheduling, achieves balanced latency across all workloads.
4.6. Discussions
While Bullet shows promising improvements, we identify several operational boundaries. First, on low-compute GPUs running dense models, the limited SM budget obliges the decode phase to claim a larger ratio of SMs to saturate memory bandwidth (Figure 8(a)), slightly tempering concurrent-execution benefits. However, this caveat is immaterial for MoE models, whose lighter compute footprint preserves the full advantage (Figure 16). Second, Bullet’s SM isolation policy could be relaxed to partial sharing, yet such exploration is orthogonal to the paper’s core contribution of fine-grained spatial–temporal orchestration and can be retrofitted without architectural changes. Third, for specialized architectures like DeepSeek-MLA (DeepSeek-AI et al., 2024) where disaggregation proves fundamentally advantageous, Bullet’s co-located approach may not match dedicated solutions’ peak performance. Extending the latency model to cover new attention variants or LoRA adapters (Hu et al., 2022) is straightforward and left to future work. Bullet’s core innovations maintain broad applicability across standard LLM architectures and typical workload distributions while extensible to emerging models.
@@ 5. Related Works
LLM Serving Systems. System-level optimizations for LLM serving have been extensively explored, which can be categorized as prefill-decode co-location and disaggregation. Co-located techniques represented by chunked prefill (Agrawal et al., 2023; Zhu et al., 2025) improve batching efficiency but maintain lockstep prefill and decode phase execution that underutilizes hardware. Disaggregated systems (Zhong et al., 2024; Patel et al., 2024; Gao et al., 2024; Qin et al., 2025) eliminate inter-phase interference but incur costly KV cache migration and load imbalance among prefill-decode instances. Bullet uniquely enables intra-GPU prefill-decode execution, maximizing utilization without overheads while meeting SLOs. Recent intra-GPU sharing proposals for LLM inference remain fundamentally constrained. Nanoflow (Zhu et al., 2025) exploits kernel overlap atop chunked prefill. However, the approach’s intrinsic limitations erode gains as sequence lengths increase. MuxServe (Duan et al., 2024) and Semi-PD (Hong et al., 2025) rely on static SM allocations, requiring costly engine relaunches for reconfiguration. Drift (Cui et al., 2025) explores lock-step execution between prefill blocks and decode in gangs using predefined SM partitions. These coarse designs lack the agility required by dynamic workloads. Bullet addresses these shortcomings through adaptive fine-grained resource management, enabling optimal SM utilization across diverse workloads.
@@
Concurrent Kernel Execution. Co-executing kernels with complementary resource demands has been proven to improve GPU utilization and performance. Spatial-temporal multiplexing techniques (Chow et al., 2023; Dhakal et al., 2020; Gu et al., 2023) typically classify kernels by computational characteristics and rely on CUDA streams (Corporation, 2025) or MPS (NVIDIA, 2025) for kernel submission. However, these approaches treat the hardware scheduler as a blackbox, providing limited control (Jain et al., 2019; Olmedo et al., 2020) over actual execution. While several systems use precise kernel timing and resource provisioning (Zhang et al., 2025a; Choi et al., 2022; Strati et al., 2024) to enforce predictable scheduling, they focus on traditional DNN workloads and fail to address LLM-specific requirements, including prefill-decode phase coordination and KV cache management. Bullet distinguishes itself from previous works by fine-grained orchestration of LLM’s phases and layer-level precise resource management.
@@
Kernel Scheduling. Effective scheduling is essential for optimizing concurrent kernel execution. Latency model-based solutions (Zhang et al., 2025a; Choi et al., 2022; Strati et al., 2024) schedule kernels within latency constraints but rely on excessive offline profiling, which is unsuitable for the dynamic workloads in LLM serving. Although estimating standalone LLM inference latency is well-established (Agrawal et al., 2024; Yuan et al., 2024; Wu et al., 2024), the complex interactions between concurrent prefill and decode phases remain unstudied. Existing solutions either rely on simplified models (Hong et al., 2025) or assume contention-free execution (Cui et al., 2025). Bullet addresses this gap by developing a profile-augmented analytical model for performance estimation and refining kernel scheduling mechanisms for LLM serving.